Circular No. 7737

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001fa IN NGC 673 M. Papenkova and W. D. Li, University of California at Berkeley

(UCB), report the discovery by LOTOSS (cf. IAUC 7514) of an apparent supernova in unfiltered images taken with the Katzman Automatic Imaging Telescope (KAIT) on Oct. 18.4 (mag ∼ 16.9) and 19.4 UT (mag ∼ 16.4). The new object is located at α = 1^{h}48^{m}22^{s}.22, δ = +11^{o}31′34′′.4 (equinox 2000.0), which is 4′′.0 west and 17′′.5 north of the nucleus of NGC 673. A KAIT image taken on Oct. 13.4 showed nothing at this position (limiting mag ∼ 19.0). SN 1996bo (IAUC 6497) also occurred in NGC 673.

A. V. Filippenko and R. Chornock, UCB, report that a CCD spectrum (range 330–1000 nm) obtained on Oct. 20 UT with the 3-m Shane reflector at Lick Observatory reveals that the object is a young type-IIn supernova, with a nearly featureless, very blue continuum and weak, narrow hydrogen-Balmer emission lines. Emission lines of He II 468.6-nm and C III/N III 464-nm, characteristic of Wolf-Rayet stars, are several times stronger than Hβ. The overall spectrum is similar to that of the type-IIn supernova 1998S at very early times (Leonard et al. 2000, Ap.J. 536, 239).

SUPERNOVAE 2001ej, 2001em, 2001eq, 2001er, AND 2001ew Filippenko and Chornock also write: “Inspection of CCD spectra, ob-

tained as above on Oct. 20 UT, reveals that SN 2001em (IAUC 7722) is of type Ib or Ic (most likely Ic), perhaps a month after maximum brightness. SN 2001eq (IAUC 7728) is also probably of type Ic. SN 2001ew (IAUC 7734) is of type Ia, < 1 week after maximum light. SN 2001er (IAUC 7728) is of type Ia, 1–2 months past maximum. SN 2001ej (IAUC 7719, 7721) is of type Ib, not Ic; the He I lines are now quite strong.”

SUPERNOVA 2001ez IN PGC 17642 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2001ez (cf. IAUC 7736), obtained by M. Calkins on Oct. 19.49 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova. Adopting the NASA/IPAC Extragalactic Database recession velocity of 3872 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hα line is 10 400 km/s.

Filippenko and Chornock add: “SN 2001ez is of type II, not long after maximum brightness; however, the P-Cyg profile of Hα exhibits much stronger emission than absorption.”

2001 October 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green